A recent poll of 2300 randomly selected 18-25-year-olds revealed that 266 currently use marijuana or hashish. According to a publication, 12.7 % of 18-25-year-olds were current users of marijuana or hashish in 1997. Do the data provide sufficient evidence to conclude that the percentage of 18-25-year-olds who currently use marijuana or hashish has changed from the 1997 percentage of 12.7%? Use α=0.05 significance level. test statistic z= positive critical z score negative critical z score
Answer #1

Sample proportion = 266 / 2300 = 0.1156

Test statistics z = - p / sqrt( p (1 - p) / n)

= 0.116 - 0.127 / sqrt( 0.127 * 0.873 / 2300)

= -1.64

Positive critical Z = 1.96

Negative critical Z = -1.96

Since test statistics falls in non-rejection region, we fail to reject null hypothesis.

We were unable to transcribe this image